1. Assume the following market demand curve: Q = 13 – 0.5P. Assume marginal cost is constant at $2.

Also assume there is a dominant firm with 6 identical competitive fringe suppliers. .

The following information will be useful for this problem:

           Residual Demand = Market Demand – Fringe Supply

           Total Fringe Supply: 3 + 0.5P

a. Calculate the equilibrium price and quantity for the dominant firm and the average fringe firm.

b. Calculate the value consumer surplus (based on total output--the dominant firm + fringe).

c. Calculate the deadweight loss (based on total output--the dominant firm + fringe).

d. Calculate the profit for the dominant firm.

e. Illustrate the Demand, market price, market quantity, consumer surplus and deadweight loss(based on total output--the dominant firm + fringe) .


1
Expert's answer
2021-10-13T15:44:47-0400
"13-0.5p=3+0.5p"

"p=10"

"Q=8"

"MC=p^\/=2"

"Q^/=2\times8=16"

"CS=\frac{26-10}{2}\times8=64"

"DL=(13-8)\times\frac{10}{2}=25"

"\pi=TR-TC=10\times8-2\times8=64"